# Hybrid quantum-classical systems: statistics, entropy, microcanonical ensemble and its connection to the canonical ensemble

### v1

@@ IV.1 The limit classical marginal distribution
where the degeneracy factor $N_{i}^{\xi}$ indicates the density of internal states due to the tracing-out of the quantum subsystem. As in (43), only one Dirac delta distribution can be different from zero at each $\xi\in M_{C}$. Notice that, for trivial quantum systems of dimension 1, degeneracy is $N_{i}^{\xi}=1$ for all states $i$ and all $\xi\in M_{C}$. In this situation, the classical marginal distribution $F_{\mu,C}^{0}(\xi)$ reproduces the classical microcanonical ensemble, showing a good classical limit behavior.
@@ IV.2 Quantum distributions in the limit
On the other hand, the quantum conditional ensemble $\tilde{\rho}_{\mu,Q,\xi}^{0}$ in the $\epsilon\to 0$ limit can be simply identified from the relation (16) applied to the $\epsilon\to 0$ limit:
@@
where the fact that, at every $\xi\in M_{C}$, there exists at most one state in the set of MESs with energy exactly equal to $E$ is taken into account. The quantum conditional ensemble is similar to the quantum microcanonical ensemble with energy $E$. However, unlike the purely quantum case, the quantum conditional ensemble for a generic hybrid system may be well defined (i.e. different from zero for some $\xi\in M_{C}$) for a continuous range of energies. In particular, if the Hamiltonian incorporates a kinetic term such as in (40), then for every energy above the minimum the quantum conditional ensemble is non zero for at least one $\xi\in M_{C}$. This new property is shared by the quantum marginal distribution in the $\epsilon\to 0$ limit
@@
with $Z_{0}$ and $\sigma_{E}$ defined in (44). For realistic hybrid systems, i.e those with unbounded kinetic or potential terms in the Hamiltonian as in (40), the quantum marginal distribution $\hat{\rho}_{\mu,Q}^{0}$ is a convex combination of $Z_{0}$ rank-one projectors $\pi_{i,j}^{\xi}$ for every energy above the minimum. In the trivial case that eigenvalues and eigenstates do not vary with $\xi$, the quantum marginal distribution becomes the microcanonical ensemble of a purely quantum system with energy $E$, which as usual is only defined if $E$ is an eigenvalue of the Hamiltonian.
V The relation between hybrid microcanonical and canonical ensembles
@@ V.1 MESs for the compound hybrid system
Consider the respective Hamiltonians of the two subsystems to the following mathematical objects: a hybrid observable $\hat{H}^{S}\in\hat{\mathcal{O}}_{H}$ for the system S, and a purely quantum observable $H^{R}\in\mathcal{O}_{Q}$ for the reservoir R, with the following eigenvalue equations:
$\hat{H}^{S}(\xi)|u_{i,j}^{S\xi}\rangle=e_{i}^{S\xi}|u_{i,j}^{S\xi}\rangle,\ H^{R}|v_{m,n}^{R}\rangle=e_{m}^{R}|v_{m,n}^{R}\rangle$ (48)
The eigenvectors, assumed to be normalized, conform an adiabatic bases $b^{S}(\xi)=\{|u_{i,j}^{S\xi}\rangle\}$ for the hybrid system S and a basis $b^{R}=\{|v_{m,n}^{R}\rangle\}$ for $\mathcal{H}^{R}$. Thus, the set $b(\xi)=\{|u_{i,j}^{S\xi}\rangle\otimes|v_{m,n}^{R}\rangle\rangle\}$ is an adiabatic basis for the whole hybrid system S+R. If a weak coupling between the system S and the reservoir R is assumed, then each state described by $|u_{i,j}^{S\xi},v_{m,n}^{R}\rangle=|u_{i,j}^{S\xi}\rangle\otimes|v_{m,n}^{R}\rangle$ has energy $e_{i}^{S\xi}+e_{m}^{R}$. Thus, a maximal set of MESs for the compound hybrid system with known energy, in the fashion of (25), can be written as follows:
@@
$\qquad\left.\pi^{\xi}_{k}=|u_{i,j}^{S\xi},v_{m,n}^{R}\rangle\langle u_{i,j}^{S\xi},v_{m,n}^{R}|\right\}$
@@ V.2 Microcanonical ensemble of the compound system
The microcanonical ensemble of the compound system is described by restricting the energy of states in $\mathcal{B}_{S+R}$ to the energy interval $E,E+\epsilon]$, obtaining as in (26) the following subset:
@@
As long as $\mathcal{A}_{S+R}^{E,\epsilon}\neq\emptyset$, the microcanonical ensemble of the compound system is given by (35):
$\hat{\rho}_{\mu}^{S+R}(\xi)=\frac{1}{\Omega}\left\{\begin{aligned} &\sum_{k\in\iota(\xi)}\pi_{k}^{\xi},&&\iota(\xi)\neq\emptyset\\ &0,&&\iota(\xi)=\emptyset\end{aligned}\right.,\quad\xi\in M_{C}^{S}$ (51)
where the normalization constant $\Omega$ represents the number of states of the compound system S+R whose energy is within the interval $E,E+\epsilon]$, i.e. the number of states in $\mathcal{A}_{S+R}^{E,\epsilon}$.
@@ V.3 Connection between the canonical and microcanonical ensembles
Once that the full compound system S+R is described, our interest is to analyze the resulting ensemble for the hybrid system S. In order to do so, we will work under the assumption that the reservoir is large enough so that, for any $\xi\in M_{C}^{S}$ such that $\iota(\xi)\neq\emptyset$, there exists at least one state of the compound system with energy $e^{T}=e^{S}+e^{R}\in[E,E+\epsilon]$. The marginal hybrid density matrix can thus be obtained by tracing out the quantum degrees of freedom corresponding to the reservoir R [18], which can be computed directly as the projectors of the compound system S+R are separable:
@@
Assuming $\epsilon$ small enough, each $\Omega^{R}_{i}$ is approximately equal to the number of states of the reservoir R with energy exactly equal to $E-e^{S,\xi}_{i}$. The number of states of a system with a specific value of energy can be computed as the exponential of the entropy of a microcanonical ensemble at said energy (taking the Boltzmann constant equal to 1). Thus, using the von Neumann’s entropy for a microcanonical ensemble $\hat{\rho}_{\mu}^{R}$ of the quantum reservoir R at energy $e_{R}$, written as $S_{N}^{R}(e_{R})=S_{N}[\hat{\rho}_{\mu}^{R}]$, the number of states with said energy is $\exp\left(S^{R}_{N}(e_{R})\right).$ If we further assume that only a small part of the total energy corresponds to the system S, i.e. $e^{S,\xi}_{i}\ll E$ for every $i$ and every $\xi\in M_{C}$, we can consider the following power expansion with respect to each $e^{S,\xi}_{i}$ truncating to first order:
@@
$\Omega^{R}_{i}\simeq\exp\left(S_{N}^{R}(E-e^{S,\xi}_{i})\right)\simeq\exp\left(S_{N}^{R}(E)\right)\exp\left(-\beta e^{S,\xi}_{i}\right)$ (54)
@@
$=\exp\left(S_{N}^{R}(E)\right)\int_{M_{C}}d\mu_{C}\,\mathrm{Tr}\left(\exp\left(-\beta\hat{H}^{S}(\xi)\right)\right)$ (55)
with $N^{\xi}_{i}$ the degeneracy at $\xi\in M_{C}$ of the $i$-th eigenvalue of the Hamiltonian $\hat{H}^{S}(\xi)$.
Substituting in (52), the marginal ensemble for system S can be obtained. Additionally, under the $e^{S,\xi}_{i}\ll E$ assumption for every $i$, every state of system S can be populated, so that the sum in (52) extends over all possible states $i,j$. In conclusion, the marginal ensemble of system S is given, up to first order, by the following hybrid density matrix:
@@
This is, as expected, the hybrid canonical ensemble for a hybrid system in contact with a reservoir, as first introduced in [13]. Within this interpretation, the constant $\beta$, introduced in (53) without any physical interpretation at the moment, can now be identified as the inverse of the temperature $T=\beta^{-1}$ of the reservoir. The partition function $Z$ is derived from the approximation (53), and allows for the normalization condition (19) to be satisfied.
The results show the connection between the hybrid microcanonical and canonical ensembles, a required property of the proposed ensemble that supports is validity as its extension to hybrid systems of the standard purely quantum and classical ensembles. It is immediate to show with similar arguments that, if considering either classical or hybrid reservoirs instead of quantum, one also obtains the hybrid canonical ensemble as the marginal distribution of system S, the only difference being the computation of the partial trace, which in every case provides as a factor the number of states with energy $e^{T}-e^{S,\xi}_{i}$.
@@ VI Microcanonical ensemble of a hybrid qubit
Let us consider a simple interaction model [45] consisting of a 2-level quantum system ($\mathcal{H}=\mathbb{C}^{2}$) coupled to a classical bath with a single parameter $R\in\mathbb{R}=M_{C}$, with hybrid Hamiltonian given by the following expression:
$\hat{H}(R)=V(R)\mathbb{I}+H_{Q}+\hat{H}_{int}(R),$ (57)
where $V(R)$ represents the potential of the classical bath, $\mathbb{I}$ is the identity operator on $\mathcal{H}$, $H_{Q}$ is the Hamiltonian of the quantum subsystem and $\hat{H}_{int}(R)$ describes the classical-quantum interaction.
Let us consider a simple interaction model [45]. In order to obtain simple but non-trivial results, we will use a quadratic bath potential and a simple quantum Hamiltonian $H_{Q}$ with eigenvalues 0 and $0.1$ in natural units ($\hbar=1$); their expressions in a basis of eigenvectors of $H_{Q}$ are:
$V(R)=R^{2},\ H_{Q}=\begin{pmatrix}0&0\\ 0&0.1\end{pmatrix},\ \hat{H}_{int}(R)=R\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ (58)
@@
$\begin{gathered}\hat{H}(0)|u^{0}_{\pm}\rangle=e^{0}_{\pm}|u^{0}_{\pm}\rangle;\quad e^{0}_{-}=0,\ e^{0}_{+}=\frac{1}{10};\\ |u^{0}_{-}\rangle=\begin{pmatrix}1\\ 0\end{pmatrix},\ |u^{0}_{+}\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}\end{gathered}$ (60)
Both eigenvalues are represented with respect to the classical parameter $R$ in Figure 2. These quantities represent the energy of the whole hybrid system when its state is $(R,|u^{R}_{-}\rangle\langle u^{R}_{-}|)$ and $(R,|u^{R}_{+}\rangle\langle u^{R}_{+}|)$, respectively. Therefore, the eigenvalues give us the necessary information in order to define the hybrid microcanonical ensemble at different energies.
Figure 2: Eigenvalues of the hybrid Hamiltonian $\hat{H}(R)$ are represented with respect to the classical parameter $R$. Additionally, energy bands of width $\epsilon=0.01$ are represented horizontally, with bottom energies equal to 0.25-0.25, 0.1-0.1, $0.05$, $0.4$. Each intersection of the energy bands with the curves of the eigenvalues is projected vertically onto the $R$ axis, defining an interval of values, as shown along the axis. These intervals have been computed numerically from (59).
@@
    Notice that the normalization factor $0.026$ is included in both non-zero pieces of the ensemble and has to be computed considering all the values of $R$, as it is necessary for the normalization of $\hat{\rho}_{\mu}(R)$ by (19).
As a last remark, notice that, in connection to what was exposed in section IV, the $\epsilon\to 0$ limit of all microcanonical ensembles described in the example are well defined and share the same properties as a classical microcanonical ensemble. For example, consider the case $E=0.05$. If the $\epsilon\to 0$ limit is taken in ensemble (62), one simply gets
@@ VII Conclusions and future work
One of the most remarkable properties of the hybrid microcanonical ensemble here presented is its ability to propagate classical properties to its quantum components. Typically, classical microcanonical ensembles can be defined for any fixed energy of the system above its minimum; the same is not true purely quantum systems, as a microcanonical ensemble can only be formulated for energies equal to the eigenvalues of the Hamiltonian, whose spectrum is generally discrete. We show in this article how, given a hybrid microcanonical ensemble, its marginal quantum ensemble is similar to a purely quantum microcanonical ensemble but, due to the influence of the classical subsystem, can be generally defined for a continuum range of energies. This extension of classical properties to quantum statistics may be useful in the analysis of more general systems than those proposed in this article, such as molecules and quantum gravity models.
We also show how the hybrid microcanonical ensemble allows us to compute the hybrid canonical ensemble proposed in [13]. By considering the microcanonical ensemble of a compound system with a hybrid subsystem and a quantum reservoir, and using the typical assumptions in this setting [18], the hybrid canonical ensemble is obtained as the marginal ensemble of the hybrid subsystem. This is another proof of the validity of the proposed hybrid microcanonical ensemble in the context of building a self-coherent statistical description of hybrid quantum-classical systems.
@@
Our work can be related with other existing approaches to hybrid systems. For example, in [15] the author focus their attention on aspects of hybrid quantum-classical ensembles which do not require introducing the entropy. However, an approach based on configuration space ensembles overcomes many difficulties arising in other approaches, as explained in section 8.1 of said reference.
Different dynamical approaches may also be considered by using new notions which generalize the idea of quantum operations (also known as quantum dynamical maps) [46] to the hybrid framework. Very recently, one possible generalization has been introduced (see Eq. (7) in [25]). We hope to expand our analysis and to include hybrid operations in our mathematical framework in the near future.

### v2

@@ IV.1 Limit classical marginal distribution
where the degeneracy factor $N_{i}^{\xi}$ indicates the density of internal states due to the tracing-out of the quantum subsystem. As in Eq. (45), only one Dirac $\delta$ distribution can be different from zero at each $\xi\in M_{C}$. Notice that, for trivial quantum systems of dimension 1, degeneracy is $N_{i}^{\xi}=1$ for all states $i$ and all $\xi\in M_{C}$. In this situation, the classical marginal distribution $F_{\mu,C}^{0}(\xi)$ reproduces the classical microcanonical ensemble, showing a good classical limit behavior.
@@ IV.2 Quantum distributions in the limit
However, the quantum conditional ensemble $\tilde{\rho}_{\mu,Q,\xi}^{0}$ in the $\epsilon\to 0$ limit can be simply identified from the relation (16) applied to the $\epsilon\to 0$ limit:
@@
where the fact that, at every $\xi\in M_{C}$, there exists at most one state in the set of MESs with energy exactly equal to $E$ is taken into account. The quantum conditional ensemble is similar to the quantum microcanonical ensemble with energy $E$. However, unlike the purely quantum case, the quantum conditional ensemble for a generic hybrid system may be well defined (i.e. different from zero for some $\xi\in M_{C}$) for a continuous range of energies. In particular, if the Hamiltonian incorporates a kinetic term such as in Eq. (42), then for every energy above the minimum the quantum conditional ensemble is non zero for at least one $\xi\in M_{C}$. This new property is shared by the quantum marginal distribution in the $\epsilon\to 0$ limit
@@
with $Z_{0}$ and $\sigma_{E}$ defined in Eq. (46). For realistic hybrid systems, i.e those with unbounded kinetic or potential terms in the Hamiltonian as in Eq. (42), the quantum marginal distribution $\hat{\rho}_{\mu,Q}^{0}$ is a convex combination of $Z_{0}$ rank-one projectors $\pi_{i,j}^{\xi}$ for every energy above the minimum. In the trivial case that eigenvalues and eigenstates do not vary with $\xi$, the quantum marginal distribution becomes the microcanonical ensemble of a purely quantum system with energy $E$, which as usual is only defined if $E$ is an eigenvalue of the Hamiltonian.
V Relation between hybrid microcanonical and canonical ensembles
@@ V.1 MESs for the compound hybrid system
Consider the respective Hamiltonians of the two subsystems: a hybrid observable $\hat{H}^{S}\in\hat{\mathcal{O}}_{H}$ for the system $S$, and a purely quantum observable $H^{R}\in\mathcal{O}_{Q}$ for the reservoir $R$, with the following eigenvalue equations:
$\hat{H}^{S}(\xi)|u_{i,j}^{S\xi}\rangle=e_{i}^{S\xi}|u_{i,j}^{S\xi}\rangle,\ H^{R}|v_{m,n}^{R}\rangle=e_{m}^{R}|v_{m,n}^{R}\rangle.$ (50)
The eigenvectors, assumed to be normalized, conform an adiabatic bases $b^{S}(\xi)=\{|u_{i,j}^{S\xi}\rangle\}$ for the hybrid system $S$ and a basis $b^{R}=\{|v_{m,n}^{R}\rangle\}$ for $\mathcal{H}^{R}$. If a weak coupling between the system $S$ and the reservoir $R$ is assumed, then each state described by $|u_{i,j}^{S\xi},v_{m,n}^{R}\rangle=|u_{i,j}^{S\xi}\rangle\otimes|v_{m,n}^{R}\rangle$ has energy $e_{i}^{S\xi}+e_{m}^{R}$. The set $b(\xi)=\{|u_{i,j}^{S\xi}\rangle\otimes|v_{m,n}^{R}\rangle\rangle\}$ is then an adiabatic basis for the whole hybrid system $S+R$. Thus, a maximal set of MESs for the compound hybrid system with known energy, in the fashion of Eq. (25), can be written as follows:
@@
$\qquad\left.\pi^{\xi}_{k}=|u_{i,j}^{S\xi},v_{m,n}^{R}\rangle\langle u_{i,j}^{S\xi},v_{m,n}^{R}|\right\}.$
The “weak coupling” condition is normally justified in terms of the size of system and reservoir: first of all, the reservoir is very large in comparison with the system (a condition that becomes important in the following step); furthermore, both system and reservoir are large enough that the frontier region between them becomes unimportant, which is what permits to neglect the interaction term.
@@ V.2 Microcanonical ensemble of the compound system
The microcanonical ensemble of the compound system is described by restricting the energy of states in $\mathcal{B}_{S+R}$ to the energy interval $E,E+\epsilon]$, obtaining as in Eq. (26) the following subset:
@@
As long as $\mathcal{A}_{S+R}^{E,\epsilon}\neq\emptyset$, the microcanonical ensemble of the compound system is given by Eq. (35):
$\hat{\rho}_{\mu}^{S+R}(\xi)=\frac{1}{\Omega}\left\{\begin{aligned} &\sum_{k\in\iota(\xi)}\pi_{k}^{\xi},&&\iota(\xi)\neq\emptyset\\ &0,&&\iota(\xi)=\emptyset\end{aligned}\right.,\quad\xi\in M_{C}^{S},$ (53)
where the normalization factor $\Omega$, computed in Eq. (36), represents the number of states of the compound system $S+R$ whose energy is within the interval $E,E+\epsilon]$, i.e. the number of states in $\mathcal{A}_{S+R}^{E,\epsilon}$.
@@ V.3 Connection between the canonical and microcanonical ensembles
Once that the full compound system $S+R$ is described, our interest is to analyze the resulting ensemble for the hybrid system $S$. To do so, we will work under the assumption that the reservoir is large enough so that, for any $\xi\in M_{C}^{S}$ such that $\iota(\xi)\neq\emptyset$, there exists at least one state of the compound system with energy $e^{T}=e^{S}+e^{R}\in[E,E+\epsilon]$. The marginal hybrid density matrix can thus be obtained by tracing out the quantum degrees of freedom corresponding to the reservoir $R$ [19], which can be computed directly as the projectors of the compound system $S+R$ are separable:
@@
Assuming $\epsilon$ small enough, each $\Omega^{R}_{i}$ is approximately equal to the number of states of the reservoir $R$ with energy exactly equal to $E-e^{S,\xi}_{i}$, as shown in Eq. (36). The number of states of a system with a specific value of energy can be computed as the exponential of the entropy of a microcanonical ensemble at said energy (taking the Boltzmann constant equal to 1). Thus, using the von Neumann’s entropy for a microcanonical ensemble $\hat{\rho}_{\mu}^{R}$ of the quantum reservoir $R$ at energy $e_{R}$, written as $S_{N}^{R}(e_{R})=S_{N}[\hat{\rho}_{\mu}^{R}]$, the number of states with said energy is $\exp\left(S^{R}_{N}(e_{R})\right).$ If we further assume that only a small part of the total energy corresponds to the system $S$, i.e. $e^{S,\xi}_{i}\ll E$ for every $i$ and every $\xi\in M_{C}$, then we can consider the following power expansion with respect to each $e^{S,\xi}_{i}$ truncating to first order:
@@
$\Omega^{R}_{i}\simeq\exp\left(S_{N}^{R}(E-e^{S,\xi}_{i})\right)\simeq\exp\left(S_{N}^{R}(E)\right)\exp\left(-\beta e^{S,\xi}_{i}\right).$ (56)
@@
$=\exp\left(S_{N}^{R}(E)\right)\int_{M_{C}}d\mu_{C}\,\mathrm{Tr}\left(\exp\left(-\beta\hat{H}^{S}(\xi)\right)\right),$ (57)
with $N^{\xi}_{i}$ the degeneracy at $\xi\in M_{C}$ of the $i$th eigenvalue of the Hamiltonian $\hat{H}^{S}(\xi)$.
Substituting in Eq. (54), the marginal ensemble for system $S$ can be obtained. Additionally, under the $e^{S,\xi}_{i}\ll E$ assumption for every $i$, every state of system $S$ can be populated, so that the sum in Eq. (54) extends over all possible states $i,j$. In conclusion, the marginal ensemble of system $S$ is given, up to first order, by the following hybrid density matrix:
@@
This is, as expected, the hybrid canonical ensemble for a hybrid system in contact with a reservoir, as first introduced in Ref. [14]. Within this interpretation, the constant $\beta$, introduced in Eq. (55) without any physical interpretation at the moment, can now be identified as the inverse of the temperature $T=\beta^{-1}$ of the reservoir. The partition function $Z$ is derived from the approximation (55), and allows for the normalization condition (19) to be satisfied.
The results show the connection between the hybrid microcanonical and canonical ensembles, a required property of the proposed ensemble that supports is validity as its extension to hybrid systems of the standard purely quantum and classical ensembles. It is immediate to show with similar arguments that, if considering either classical or hybrid reservoirs instead of quantum, then one also obtains the hybrid canonical ensemble as the marginal distribution of system $S$, the only difference being the computation of the partial trace, which in every case provides as a factor the number of states with energy $e^{T}-e^{S,\xi}_{i}$.
@@ VI Microcanonical ensemble of a hybrid qubit
Let us consider a simple interaction model consisting of a two-level quantum system ($\mathcal{H}=\mathbb{C}^{2}$) coupled to a classical bath with a single parameter $R\in\mathbb{R}=M_{C}$, with hybrid Hamiltonian given by the following expression:
$\hat{H}(R)=V(R)\mathbb{I}+H_{Q}+\hat{H}_{\rm int}(R),$ (59)
where $V(R)$ represents the potential of the classical bath, $\mathbb{I}$ is the identity operator on $\mathcal{H}$, $H_{Q}$ is the Hamiltonian of the quantum subsystem and $\hat{H}_{\rm int}(R)$ describes the classical-quantum interaction. The model has been used [50] (see also Ref. [51]) as an effective model for a dissipative quantum system where only two energy levels have physical relevance (although it can be generalized to higher finite dimensional cases). Notice that, from a hybrid system point of view, it is a particularly simple case, since classical degrees of freedom are not dynamical for they are acting as a bath for the quantum ones. Nonetheless, our MaxEnt construction is not dependent on the system dynamics, and therefore we decided to choose this simple yet interesting model to exemplify our construction without computational difficulties.
To obtain simple but nontrivial results, we will use a quadratic bath potential and a simple quantum Hamiltonian $H_{Q}$ with eigenvalues 0 and $0.1$ in natural units ($\hbar=1$); their expressions in a basis of eigenvectors of $H_{Q}$ are
$V(R)=R^{2},\ H_{Q}=\begin{pmatrix}0&0\\ 0&0.1\end{pmatrix},\ \hat{H}_{\rm int}(R)=R\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. (60)
@@
$\begin{gathered}\hat{H}(0)|u^{0}_{\pm}\rangle=e^{0}_{\pm}|u^{0}_{\pm}\rangle;\quad e^{0}_{-}=0,\ e^{0}_{+}=\frac{1}{10};\\ |u^{0}_{-}\rangle=\begin{pmatrix}1\\ 0\end{pmatrix},\ |u^{0}_{+}\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}\end{gathered}$. (62)
Both eigenvalues are represented with respect to the classical parameter $R$ in Fig. 2. These quantities represent the energy of the whole hybrid system when its state is $(R,|u^{R}_{-}\rangle\langle u^{R}_{-}|)$ and $(R,|u^{R}_{+}\rangle\langle u^{R}_{+}|)$, respectively. Therefore, the eigenvalues give us the necessary information to define the hybrid microcanonical ensemble at different energies.
Figure 2: Eigenvalues $e_{+}^{R}$ and $e_{-}^{R}$ of the hybrid Hamiltonian $\hat{H}(R)$, shown in Eqs. (61) and (62), are represented as functions of the classical parameter $R$. Additionally, energy bands of width $\epsilon=0.01$ are represented horizontally, with bottom energies equal to 0.25-0.25, 0.1-0.1, $0.05$, $0.4$. Each intersection of the energy bands with the curves of the eigenvalues is projected vertically onto the $R$ axis, defining an interval of values, as shown along the axis. These intervals have been computed numerically from Eq. (61).
@@
    Notice that the normalization factor $0.026$ is included in both nonzero pieces of the ensemble and has to be computed considering all the values of $R$, as it is necessary for the normalization of $\hat{\rho}_{\mu}(R)$ by Eq. (19).
As a last remark, notice that, in connection to what was exposed in Sec. IV, the $\epsilon\to 0$ limit of all microcanonical ensembles described in the example are well defined and share the same properties as a classical microcanonical ensemble. For example, consider the case $E=0.05$. If the $\epsilon\to 0$ limit is taken in ensemble (64), then one simply gets
@@ VII Conclusions and future work
One of the most remarkable properties of the hybrid microcanonical ensemble here presented is its ability to propagate classical properties to its quantum components. Typically, classical microcanonical ensembles can be defined for any fixed energy of the system above its minimum. The same is not true purely quantum systems, as a microcanonical ensemble can only be strictly formulated for energies equal to the eigenvalues of the Hamiltonian, whose spectrum is generally discrete; although, naturally, for large systems any feasible small range of energies always contains some eigenvalues and the discreteness of the spectrum is not a problem in practice. We show in this article how, given a hybrid microcanonical ensemble, its marginal quantum ensemble is similar to a purely quantum microcanonical ensemble but, due to the influence of the classical subsystem, can be generally defined for a continuum range of energies. This extension of classical properties to quantum statistics may be key to the description of a framework in which the microcanonical ensemble can be intrinsically applied to any system, without the necessity for additional properties such as large systems, and for this reason may be useful in the analysis of more general systems than those proposed in this article, such as molecules and quantum gravity models.
We also show how the hybrid microcanonical ensemble allows us to compute the hybrid canonical ensemble proposed in Ref. [14]. By considering the microcanonical ensemble of a compound system with a hybrid subsystem and a quantum reservoir, and using the typical assumptions in this setting [19], the hybrid canonical ensemble is obtained as the marginal ensemble of the hybrid subsystem. This is another proof of the validity of the proposed hybrid microcanonical ensemble in the context of building a self-coherent statistical description of hybrid quantum-classical systems.
@@
Our work can be related with other existing approaches to hybrid systems. For example, in Ref. [16] the author focus their attention on aspects of hybrid quantum-classical ensembles which do not require introducing the entropy. However, an approach based on configuration space ensembles overcomes many difficulties arising in other approaches, as explained in Sec. 8.1 of said reference.
Different dynamical approaches may also be considered by using new notions which generalize the idea of quantum operations (also known as quantum dynamical maps [52]) to the hybrid framework. Very recently, one possible generalization has been introduced (see Eq. (7) in Ref. [28]). Approaches based on a hybrid generalization of Koopman classical dynamics [41, 53, 54, 55], or nonlinear approaches based on the restriction of Lioville Ehrenfest dynamics to a finite number of quantum moments [48] may offer interesting candidates for hybrid equilibrium dynamics. Also, recent works deal with the description of out-of-equilibrium thermodynamics of hybrid quantum-classical systems [56, 57], which provide a fertile ground for the application of entropy-based analysis of hybrid systems. Other relevant works deal with different aspects of dynamics of hybrid systems, such as path integrals [59, 58]. We hope to expand our analysis and develop our formalism to incorporate all of these cutting-edge developments in a complete mathematical framework.